A machine of force P pulls a treasure chest along a 8 foot inclined surface with an angle of elevation 20 degree. Assume the treasure chest has 40kg in mass and force P is 5 kiloNewtons applied parallel to surface , please solve:

a) work due to gravity

b) Work due to the external force

c) work due to friction if knetic friction coeff = 0.25
Answer #1

(a)Work done by force,

W = F . d (dot product)

Work done by gravity,

Wg= Fg . d =Fg* d * cos(theta)

Angle btween gravotational force and displacement,

theta = -(90-20) = 70

Wg = m* g* d* cos (-70)

= -40*9.81*8*0.3048*0.3420

= -327.23 J

(b)

Work due to external force,

W= F. d

= 5*10^3*8*.3048

=12192 J

(c)

Work due to kinetic friction force,

Wf =-Ff. d

=-mu*m*g*d

=-0.25*40*9.81*8*.3048

=-239.2 J